OpenStudy (anonymous):

Identify the Exponent Law shown in the example below. x2x4 = x2+4 = x6 1.None of the choices below are correct. 2.Power of a Product 3.Power of a Power 4.Product of Powers

OpenStudy (azureilai):

http://www.mathsisfun.com/algebra/exponent-laws.html

OpenStudy (mathstudent55):

This is what these three rules mean. Can you figure out which one was used? Power of a Product: \(\large (ab)^m = a^mb^m\) Power of a Power: \(\large (a^m)^n = a^{mn} \) Product of Powers: \(\large a^m \times a^n = a^{m+n} \)

OpenStudy (anonymous):

yeah thank you

OpenStudy (mathstudent55):

You're welcome.